# Discovery and Preliminary Characterization of a Third Interstellar Object: 3I/ATLAS

### v2

@@ Abstract
We report initial observations aimed at the characterization of a third interstellar object candidate. This object, 3I/ATLAS or C/2025 N1 (ATLAS), was discovered on 2025 July 1 UT and has an orbital eccentricity of $e\sim 6.1$, perihelion of $q\sim 1.36$ au, inclination of $\sim 175^{\circ}$, and hyperbolic velocity of $V_{\infty}\sim 58$ km s-1. We report deep stacked images obtained using the Canada-France-Hawaii Telescope and the Very Large Telescope that resolve a compact coma. Using images obtained from several smaller ground-based telescopes, we find minimal light curve variation for the object over a $\sim 4$ day time span. The visible/near-infrared spectral slope of the object is 17.1$\pm$0.2 %/100 nm, comparable to other interstellar objects and primitive solar system small bodies (comets and D-type asteroids), although this result is likely affected by some coma contamination. 3I/ATLAS will be observable through early September 2025, then unobservable by Earth-based observatories near perihelion due to low solar elongation. It will be observable again from the ground in late November 2025. Although this limitation unfortunately prohibits detailed Earth-based observations at perihelion when the activity of 3I/ATLAS is likely to peak, spacecraft at Mars could be used to make valuable observations at this time. Additional photometric, spectroscopic, and polarimetric monitoring of 3I/ATLAS by ground- and space-based telescopes, and possibly spacecraft based at Mars, are highly encouraged for characterizing 3I/ATLAS’s rotational light curve, activity evolution, nongravitational acceleration, and compositional indicators of formation conditions.
@@ 1 Introduction
These first interstellar objects displayed divergent properties. For one, 1I/‘Oumuamua displayed no visible activity (K. J. Meech et al., 2017; Q.-Z. Ye et al., 2017; D. Jewitt et al., 2017; D. E. Trilling et al., 2018), yet nongravitational acceleration was detected in its trajectory at the $\sim 30\sigma$ level (M. Micheli et al., 2018). Meanwhile, 2I/Borisov displayed clear outgassing and dust ejection activity (D. Jewitt & J. Luu, 2019; A. Fitzsimmons et al., 2019; B. T. Bolin et al., 2020; Q. Ye et al., 2020; A. J. McKay et al., 2020; P. Guzik et al., 2020; M.-T. Hui et al., 2020; Y. Kim et al., 2020; G. Cremonese et al., 2020; B. Yang et al., 2021). The excess velocity of 1I/‘Oumuamua and 2I/Borisov also differed significantly. These objects had velocities of $V_{\infty}\sim 26$ km s-1 and $V_{\infty}\sim 32$ km s-1 respectively, which approximately correspond to ages of $\sim 10^{2}$ and $\sim 10^{3}$ Myr (E. Mamajek, 2017; E. Gaidos et al., 2017; F. Feng & H. R. A. Jones, 2018; F. Almeida-Fernandes & H. J. Rocha-Pinto, 2018; T. Hallatt & P. Wiegert, 2020; C.-H. Hsieh et al., 2021). In addition, 1I/‘Oumuamua displayed brightness variations of $\sim 3.5$ magnitudes corresponding to an extreme oblate $6:6:1$ geometry (K. J. Meech et al., 2017; M. M. Knight et al., 2017; B. T. Bolin et al., 2018; W. C. Fraser et al., 2018; M. J. S. Belton et al., 2018; S. Mashchenko, 2019; A. G. Taylor et al., 2023) and had a moderately red reflectance spectrum (K. J. Meech et al., 2017; A. Fitzsimmons et al., 2018; Q.-Z. Ye et al., 2017).
@@ 2.1 Discovery
Six hours after the initial detection by ATLAS on 2025 July 1, prediscovery detections from 2025 June 28–29 were identified in Zwicky Transient Facility (ZTF; E. C. Bellm et al., 2019; M. J. Graham et al., 2019) survey data, extending the orbital arc from 3.3 hours to 3 days. The new 3-day arc suggested a strongly hyperbolic orbit and prompted speculation on community mailing lists that the object might be interstellar. Later on the early hours of 2025 July 2 UT, additional prediscovery pairs and triplets were identified in ATLAS data from five days earlier by S. Deen and later refined and resubmitted to the MPC by the ATLAS team. More prediscovery detections from ZTF, dating back to 2025 June 14222Another batch of ZTF prediscovery detections, dating back to 2025 May 22, were later submitted and published in MPEC 2025-N51 (https://minorplanetcenter.net/mpec/K25/K25N51.html)., were also identified and submitted.
@@ 2.2 Orbit Characterization
As of the early hours of 2025 July 2, the initial discovery and prediscovery arc, covering a total of about 18 days, was sufficient to confirm the large eccentricity and, consequently, hyperbolic nature of the object’s orbit. As soon as the unusual nature of the object became evident, a large number of follow-up observations were obtained on 2025 July 2 by various observatories, leading to additional orbit refinement.
Within our team, we first acquired 31 unfiltered 163 s exposures starting at 06:34 2025 UT July 2 spanning a total of 93 minutes using the European Space Agency’s (ESA) 0.56-meter Test Bed Telescope (TBT; MPC code W57) at La Silla Observatory in Chile, which has a field of view of 2.5 $\times$ 2.5. The telescope is fully devoted to NEO survey and follow-up observations, and can be interrupted at any time for high-profile targets.
@@
A successful DDT request to the European Southern Observatory’s (ESO) 8.2-meter Very Large Telescope (VLT), submitted on the same day and quickly approved, also allowed our team to obtain good-seeing images of the object on 2025 July 4. High-precision astrometry (better than ±$\pm 0\farcs 1$) was extracted from this dataset, using an extrapolation to zero aperture to ensure correction for possible asymmetries in the inner coma of the object.
Our best estimate of the heliocentric orbital elements of the object, at the time of submission of this paper, is listed in Table 1. In Figure 2, we show the orbit of the object in comparison to previously discovered interstellar objects.
Table 1: Initial orbit of 3I/ATLAS, computed using astrometry from 2025 May 22 to 2025 July 6. Heliocentric orbital elements at the epoch of 2025 July 3 TDB.
@@
3I/ATLAS is predicted to approach within 0.19 au of Mars on 2025 October 3.
@@ 3.1 Activity
Given the stark difference in visible activity levels between 1I/‘Oumuamua and 2I/Borisov (Section 1), a high priority for initial observational analysis for 3I/ATLAS was determining whether it was active or not. With this in mind, we obtained three 60 s non-sidereally guided gri-band images with the MegaCam wide-field mosaic imager (O. Boulade et al., 2003) on the 3.6-meter Canada-France-Hawaii Telescope (CFHT, MPC code T14) on 2025 July 2 to search for faint cometary activity. The highest-quality image had typical stellar FHWMs of $0\farcs 72\pm 0\farcs 05$ measured perpendicular to the direction of trailing, while the object had a FWHM of $1\farcs 29\pm 0\farcs 02$. The magnitude in a $5\farcs$ radius aperture was 17.2 in the Gaia DR2 G band after three background stars were masked, although the field was crowded, so we consider this measurement to be unreliable. Figure 3a shows the stacked composite image of these data in which faint activity is visible.
On 2025 July 4, fifteen 10 s R-band images were acquired using the FORS2 instrument on the 8.2-meter Unit Telescope 1 of the European Southern Observatory Very Large Telescope (VLT) on Paranal, Chile. The measured image quality was $0\farcs 6$ FWHM. However, the object had a FWHM of $1\farcs 49\times 1\farcs 29$ with the asymmetric elongation pointing to position angle $277^{\circ}$ from North to East, where the coma can be seen out to $3\farcs 5$. A composite image of these data is shown in Figure 3b. The photometric profile of the object (Figure 4) shows an excess extending over $4\farcs$ compared to that of field stars.
@@
Figure 5: Absolute magnitudes computed from ZTF precovery data published in MPEC 2025-N51 from 2025 May 22 to June 18. The median $H_{V}$ and standard deviation of data are shown by a dashed horizontal line and dotted horizontal lines, respectively. No photometric uncertainties were reported for these data.
As noted in Section 2, $g^{\prime}$- and $r^{\prime}$-band precovery data from ZTF were published in MPEC 2025-N51 following these observations. This included photometry dating back to 2025 May 22. These data mostly consisted of single $g^{\prime}$- or $r^{\prime}$-band detections on each night, but included pairs of $g^{\prime}$ and $r^{\prime}$-band detections on 2025 May 22 and June 18. We compute inferred $V$-band absolute magnitude from these data assuming a standard asteroidal phase function (E. Bowell et al., 1989) assuming $G=0.15$ and solar colors (K. Jordi et al., 2006; J. Holmberg et al., 2006). We use this to present a secular light curve in Figure 5. The computed absolute magnitudes are relatively constant between 2025 May 22 and June 14 (during which the object ranged between heliocentric distances of $r=5.8$ au to $r=5.1$ au). The secular light curve has a median value of $H_{V}=12.4$ and a standard deviation of 0.07 mag. The photometry from 2025 June 18 (the last night plotted in Figure 5 when the object was at $r=4.9$ au) corresponds to an absolute magnitude outside that standard deviation at $H_{V}=12.1$. For comparison, $r^{\prime}$-band light curve data obtained from LCO-network telescopes (discussed in more detail in Section 3.2) show an approximate average $r^{\prime}$-band apparent magnitude of $m_{r}\sim 17.9$, corresponding to $H_{V}=11.9$, indicating that the faint coma visible in CFHT and VLT data obtained on 2025 July 2 and 4 could be responsible for about 0.5 mag of excess brightening of the object since early June.
We note that this finding does not exclude the presence of coma prior to mid-June, but indicates that any coma that may have been present was relatively steady-state from late May to early June, and then increased in mid-June. For reference, $H_{V}=12.4$ corresponds to a radius of 10–11 km for an albedo of 0.04–0.05 typical of comets, but this is an upper limit given that we cannot rule out activity during the observations used to derive this result.
Figure 6: Compiled light curve of 3I/ATLAS incorporating $r^{\prime}$-band data from the LCO 0.35-meter telescopes, Faulkes Telescope North, Faulkes Telescope South, and the Telescope Joan Oró, Johnson-Cousins $Rc$-band data from the TRAPPIST-North and -South telescopes, and $o$-band data from ATLAS-HKO and ATLAS-CHL.
@@ 3.2 Rotation
Photometric observations were obtained on the nights of 2025 July 2–4 with the TRAPPIST telescopes (E. Jehin et al., 2011). TRAPPIST-South (TS, MPC code I40) is located at the ESO La Silla Observatory in Chile, and TRAPPIST-North (TN, MPC code Z53) is located at the Oukaimeden observatory in Morocco. Both telescopes are 0.6-meter Ritchey-Chretien telescopes operating at f/8. TS is equipped with a FLI ProLine 3041-BB CCD camera with a $22\farcm$ field of view and a pixel scale of $0\farcs 64$ pixel-1. The TN camera is an Andor IKONL BEX2 DD camera providing a $20\farcm$ field of view and pixel scale of $0\farcs 60$ pixel-1. Johnson-Cousin Rc filters were used with $2\times 2$ binning ($1\farcs 2$ pixel-1) and an exposure time of 180 s. Photometry was performed with the PhotometryPipeline (M. Mommert, 2017a) by matching $\sim$100 field stars in each image with the Pan-STARRS DR1 photometric catalog and using an aperture radius of 4 pixels.
@@
Compiled light curve results are shown in Figure 6. Apart from the differences produced by the variety of filters used, offsets between datasets are likely due to background stars and differences in pipeline removal processes. No periodicity is immediately apparent within any of the datasets, indicating little brightness variation during the observation period. This is in notable contrast to 1I/‘Oumuamua, which displayed a far more extreme light curve range of $\sim$ 3.5 magnitudes (see Section 1). It should be noted, however, that the detection of activity and coma (Section 3.1) means that any photometric variations caused by a rotating nucleus will be at least partially suppressed by the relatively more steady state coma surrounding it.
We note that accurate photometry of much of our time series data was impacted by background source interference as 3I/ATLAS passes through the Galactic plane. Additional follow-up observations of 3I/ATLAS once it leaves the Galactic plane region of the sky and its background fields become less crowded will constrain the rotation period.
@@ 3.3 Colors
In Figure 7, we show the resulting four-color surface reflectance spectrum of the object from the FTN.
Figure 7: The $g^{\prime}$,$r^{\prime}$,$i^{\prime}$,$z^{\prime}$ colors of 3I/ATLAS obtained with FTN as well as the reflectance spectrum obtained with SNIFS on the UH 2.2-meter at Maunakea normalized at 5500 $\mathrm{\AA}$ and plotted against 1I/‘Oumuamua (Q.-Z. Ye et al., 2017), 2I/Borisov (J. de León et al., 2020), the extremely red Centaur Pholus (R. P. Binzel, 1992), and the mean D-type asteroid spectrum (F. E. DeMeo et al., 2009). Note that errors on the color measurement are plotted, but are approximately the size of the plot points. The region of the spectrum masked due to the SNIFS dichroic sensitivity is highlighted in red. Both the observed spectrum and the four-color derived reflectance slopes are in agreement. 3I/ATLAS shows a moderately red spectral slope similar to 1I/‘Oumuamua and 2I/Borisov.
The colors of 3I/ATLAS as seen in Figure 7 are relatively linear (e.g., without obvious absorption features or spectral curvature) and significantly redder than those of the Sun. This reflectance spectrum is approximately 18 %/100 nm as derived from the full range of the available color data. It is similar, though somewhat redder, to that retrieved for 1I/‘Oumuamua (10$\pm$6 %/100 nm, Q.-Z. Ye et al., 2017) and 2I/Borisov (12$\pm$1 %/100 nm J. de León et al., 2020), both of which are slightly redder than the D-type asteroids (F. E. DeMeo et al., 2009). However, it is not as red as seen in some outer Solar System objects (e.g. Pholus, R. P. Binzel, 1992). Some of the reflected light must be from the coma because 3I/ATLAS is weakly active. This is corroborated by the fact that photometry of $g^{\prime}$- and $r^{\prime}$-band ZTF precovery data (Section 3.1) obtained when the object appeared to be less active, indicated that the $g^{\prime}-r^{\prime}$ colors on those nights were extremely close to solar. Specifically, the measurements were $(g^{\prime}-r^{\prime})=0.42$ on 2025 May 22 and $(g^{\prime}-r^{\prime})=0.44$ on June 18, compared to $(g^{\prime}-r^{\prime})_{\odot}=0.45$ (J. Holmberg et al., 2006). This suggests that the much redder color measured from later data could be heavily affected by ejected dust. The spectral slope measured here is slightly higher than the average for cometary dust (see modeling and discussions of typical Solar System comets in S. Protopapa et al. 2018; T. Kareta et al. 2023, as well as Figure 3 of L. Kolokolova et al. 2024), but not significantly so. 2I/Borisov’s coma was similarly red (see, e.g., J. de León et al. 2020), but the object was also significantly more active at the time of the color observations.

### v3

@@ Abstract
We report initial observations aimed at the characterization of a third interstellar object. This object, 3I/ATLAS or C/2025 N1 (ATLAS), was discovered on 2025 July 1 UT and has an orbital eccentricity of $e\sim 6.1$, perihelion of $q\sim 1.36$ au, inclination of $\sim 175^{\circ}$, and hyperbolic velocity of $V_{\infty}\sim 58$ km s-1. We report deep stacked images obtained using the Canada-France-Hawaii Telescope and the Very Large Telescope that resolve a compact coma. Using images obtained from several smaller ground-based telescopes, we find minimal light curve variation for the object over a $\sim 4$ day time span. The visible/near-infrared spectral slope of the object is 17.1$\pm$0.2 %/100 nm, comparable to other interstellar objects and primitive solar system small bodies (comets and D-type asteroids). 3I/ATLAS will be observable through early September 2025, then unobservable by Earth-based observatories near perihelion due to low solar elongation. It will be observable again from the ground in late November 2025. Although this limitation unfortunately prohibits detailed Earth-based observations at perihelion when the activity of 3I/ATLAS is likely to peak, spacecraft at Mars could be used to make valuable observations at this time.
@@ 1 Introduction
These first interstellar objects displayed divergent properties. For one, 1I/‘Oumuamua displayed no visible activity (K. J. Meech et al., 2017; Q.-Z. Ye et al., 2017; D. Jewitt et al., 2017; D. E. Trilling et al., 2018), yet nongravitational acceleration was detected in its trajectory at the $\sim 30\sigma$ level (M. Micheli et al., 2018). Meanwhile, 2I/Borisov displayed clear outgassing and dust ejection activity (D. Jewitt & J. Luu, 2019; A. Fitzsimmons et al., 2019; Q. Ye et al., 2020a; A. J. McKay et al., 2020; P. Guzik et al., 2020; M.-T. Hui et al., 2020; Y. Kim et al., 2020; G. Cremonese et al., 2020; B. Yang et al., 2021). The excess velocity of 1I/‘Oumuamua and 2I/Borisov also differed significantly. These objects had velocities of $V_{\infty}\sim 26$ km s-1 and $V_{\infty}\sim 32$ km s-1 respectively, which approximately correspond to ages of $\sim 10^{2}$ and $\sim 10^{3}$ Myr (E. Mamajek, 2017; E. Gaidos et al., 2017; F. Feng & H. R. A. Jones, 2018; F. Almeida-Fernandes & H. J. Rocha-Pinto, 2018; T. Hallatt & P. Wiegert, 2020; C.-H. Hsieh et al., 2021). In addition, 1I/‘Oumuamua displayed brightness variations of $\sim 3.5$ magnitudes corresponding to an extreme oblate $6:6:1$ geometry (K. J. Meech et al., 2017; M. M. Knight et al., 2017; W. C. Fraser et al., 2018; M. J. S. Belton et al., 2018; S. Mashchenko, 2019; A. G. Taylor et al., 2023) and had a moderately red reflectance spectrum (K. J. Meech et al., 2017; A. Fitzsimmons et al., 2018; Q.-Z. Ye et al., 2017).
@@ 2.1 Discovery
Six hours after the initial detection by ATLAS on 2025 July 1, pre-discovery detections from 2025 June 28–29 were identified in Zwicky Transient Facility (ZTF; E. C. Bellm et al., 2019; M. J. Graham et al., 2019) survey data, extending the orbital arc from 3.3 hours to 3 days. The new 3-day arc suggested a strongly hyperbolic orbit and prompted speculation on community mailing lists that the object might be interstellar. Later during the early hours of 2025 July 2 UT, additional pre-discovery pairs and triplets were identified in ATLAS data from five days earlier by S. Deen and later refined and resubmitted to the MPC by the ATLAS team. More pre-discovery detections from ZTF, dating back to 2025 June 14222Another batch of ZTF pre-discovery detections, dating back to 2025 May 22, were later submitted and published in MPEC 2025-N51 (https://minorplanetcenter.net/mpec/K25/K25N51.html)., were also identified and submitted.
@@ 2.2 Orbit Characterization
As of the early hours of 2025 July 2, the initial discovery and pre-discovery arc, covering a total of about 18 days, was sufficient to confirm the large eccentricity and, consequently, hyperbolic nature of the object’s orbit. As soon as the unusual nature of the object became evident, a large number of follow-up observations were obtained on 2025 July 2 by various observatories, leading to additional orbit refinement.
We first acquired 31 unfiltered 163 s exposures starting at 06:34 UT on 2025 July 2 spanning a total of 93 minutes using the European Space Agency’s (ESA) 0.56-meter Test Bed Telescope (TBT; MPC code W57) at La Silla Observatory in Chile, which has a field of view of 2.5 $\times$ 2.5. The telescope is fully devoted to NEO survey and follow-up observations, and can be interrupted at any time for high-profile targets.
@@
A successful DDT request to the European Southern Observatory’s (ESO) 8.2-meter Very Large Telescope (VLT), submitted on the same day and quickly approved, also allowed our team to obtain images of the object during good seeing on 2025 July 4. High-precision astrometry (better than ±$\pm 0\farcs 1$) was extracted from this dataset, using an extrapolation to zero aperture to ensure correction for possible asymmetries in the inner coma of the object.
Our best estimate of the heliocentric orbital elements of the object, at the time of submission of this paper, is listed in Table 1. In Figure 2, we show the orbit of 3I/ATLAS in comparison to previously discovered interstellar objects.
Table 1: Initial orbit of 3I/ATLAS, computed using astrometry from 2025 May 22 to 2025 July 6. Heliocentric orbital elements at the epoch of 2025 July 3.
@@ 3.1 Activity
Given the stark difference in visible activity levels between 1I/‘Oumuamua and 2I/Borisov (Section 1), a high priority for initial observational analysis for 3I/ATLAS was determining whether it was active or not. With this in mind, we obtained three 60 s non-sidereally guided gri-band images with the MegaCam wide-field mosaic imager (O. Boulade et al., 2003) on the 3.6-meter Canada-France-Hawaii Telescope (CFHT, MPC code T14) on 2025 July 2 to search for faint cometary activity. The highest-quality image had typical stellar FHWMs of $0\farcs 72\pm 0\farcs 05$ measured perpendicular to the direction of trailing, while the object had a FWHM of $1\farcs 29\pm 0\farcs 02$. The magnitude in a $5\farcs 0$ radius aperture was 17.2 in the Gaia DR2 G band after three background stars were masked, although the field was crowded, so we consider this measurement to be unreliable. Figure 3a shows the stacked composite image of these data in which faint activity is visible.
On 2025 July 4, fifteen 10 s R-band images were acquired using the FORS2 instrument on the 8.2-meter Unit Telescope 1 of the European Southern Observatory Very Large Telescope (VLT) on Cerro Paranal, Chile. The measured stellar image quality was $0\farcs 6$ FWHM. However, the object had a FWHM of $1\farcs 49\times 1\farcs 29$ with the asymmetric elongation pointing to position angle $277^{\circ}$ from North to East, where the coma can be seen extending out to $3\farcs 5$. A composite image of these data is shown in Figure 3b. The photometric profile of the object (Figure 4) shows a brightness excess extending over $4\farcs 0$ compared to that of field stars.
@@
Figure 5: Absolute magnitudes computed from ZTF precovery data published in MPEC 2025-N51 from 2025 May 22 to June 18 (black filled dots). The median $H_{V}$ and standard deviation of data are shown by a dashed horizontal line and dotted horizontal lines, respectively. No uncertainties were reported for these original data, and so plotted uncertainties are estimated based on the standard deviation of photometric points measured between 2025 May 22 and June 14. Corresponding pre-perihelion $H_{V}$ for 2I/Borisov derived from Q. Ye et al. (2020b) are plotted on the same heliocentric distance ($r$) scale (open circles) for comparison.
As noted in Section 2, $g^{\prime}$- and $r^{\prime}$-band precovery data from ZTF were published in MPEC 2025-N51 following these observations. This included photometry dating back to These data mostly consisted of single $g^{\prime}$- or $r^{\prime}$-band detections on each night, but included pairs of $g^{\prime}$ and $r^{\prime}$-band detections on 2025 May 22 and June 18. We compute inferred $V$-band absolute magnitude from these data assuming a standard asteroidal phase function (E. Bowell et al., 1989) assuming $G=0.15$ and solar colors (K. Jordi et al., 2006; J. Holmberg et al., 2006). We use this to present a secular light curve in Figure 5. The computed absolute magnitudes are relatively constant between 2025 May 22 and June 14 (during which the object ranged between heliocentric distances of $r=5.8$ au to $r=5.1$ au). The secular light curve has a median value of $H_{V}\sim 12.4$ and a standard deviation of $\sim$0.1 mag.
The photometry from 2025 June 18 (the last night plotted in Figure 5 when the object was at $r=4.9$ au) corresponds to an absolute magnitude outside that standard deviation at $H_{V}=12.1\pm 0.1$. For comparison, $r^{\prime}$-band light curve data obtained from LCO-network telescopes (discussed in more detail in Section 3.2) show an approximate average $r^{\prime}$-band apparent magnitude of $m_{r}\sim 17.9$, corresponding to $H_{V}=11.9$, indicating that the faint coma visible in CFHT and VLT data obtained on 2025 July 2 and 4 could be responsible for about 0.5 mag of excess brightening of the object since early June.
We note that this finding does not exclude the presence of coma prior to mid-June, but indicates that any coma that may have been present was relatively steady-state from late May to early June, and then increased in mid-June. For reference, $H_{V}=12.4\pm 0.1$ corresponds to a radius of $\sim(10\pm 1)$ km for a geometric albedo of $p_{V}=0.05$ typical of comets, but this is an upper limit given that we cannot rule out activity during the observations used to derive this result.
[…]
@@ 3.2 Rotation
Table 2: Light Curve Observations Summary
MPC Observation Pixel Photometry Aperture
Facility/Instrument Site Code Dates (UT) Filter(s) Scale Method Radiusa
ATLAS (Haleakala) W68 2025 July 4 $o$ $1\farcs 86$ PSF n/a
ATLAS (Chile) T05 2025 July 4 $o$ $1\farcs 86$ PSF n/a
Faulkes Telescope North F65 2025 July 2, 4 $r^{\prime}$ $0\farcs 28$ aperture $2\farcs 2$
Faulkes Telescope South E10 2025 July 2, 4 $r^{\prime}$ $0\farcs 28$ aperture $2\farcs 2$
LCO 0.35m (Haleakala #1) T03 2025 July 2, 4 $r^{\prime}$ $0\farcs 74$ aperture $3\farcs 0$
LCO 0.35m (Haleakala #2) T04 2025 July 2, 4 $r^{\prime}$ $0\farcs 74$ aperture $3\farcs 0$
Telescope Joan Oró C65 2025 July 4-5 $r^{\prime}$ $0\farcs 72$ aperture $2\farcs 2$
TRAPPIST-North Z53 2025 July 2-4 $R_{c}$ $1\farcs 20$ aperture $4\farcs 8$
TRAPPIST-South I40 2025 July 2-4 $R_{c}$ $1\farcs 28$ aperture $5\farcs 1$
a Photometry aperture size if aperture photometry is used; otherwise, this field is not applicable (n/a).
@@
Photometric observations were obtained on the nights of 2025 July 2–4 with the TRAPPIST telescopes (E. Jehin et al., 2011). TRAPPIST-North (TN, MPC code Z53) is located at the Oukaimeden observatory in Morocco, and TRAPPIST-South (TS, MPC code I40) is located at the ESO La Silla Observatory in Chile. Both telescopes are 0.6-meter Ritchey-Chretien telescopes operating at f/8. The TN camera is an Andor IKONL BEX2 DD camera providing a $20^{\prime}$ field of view and pixel scale of . TS is equipped with a FLI ProLine 3041-BB CCD camera with a $22^{\prime}$ field of view and a pixel scale of . Johnson-Cousin $R_{c}$ filters were used with individual exposure times of 180 s .
@@
Compiled light curve results are shown in Figure 6. .
No periodicity is immediately apparent within any of the datasets, indicating little brightness variation during the observation period This is in notable contrast to 1I/‘Oumuamua, which displayed a far more extreme light curve range of $\sim$ 3.5 magnitudes (see Section 1). It should be noted, however, that the detection of activity and coma (Section 3.1) means that any photometric variations caused by a rotating nucleus will be at least partially suppressed by the relatively more steady state coma surrounding it.
@@ 3.3 Colors
[…]
The colors of 3I/ATLAS as seen in Figure 7 are relatively linear (e.g., without obvious absorption features or spectral curvature) and significantly redder than those of the Sun. This slope of this reflectance spectrum is approximately 18 %/100 nm as derived from the full range of the available color data. It is similar, though somewhat redder, to that retrieved for 1I/‘Oumuamua (10$\pm$6 %/100 nm, Q.-Z. Ye et al., 2017) and 2I/Borisov (12$\pm$1 %/100 nm J. de León et al., 2020), both of which are slightly redder than the D-type asteroids (F. E. DeMeo et al., 2009). However, it is not as red as seen in some outer Solar System objects (e.g. Pholus, R. P. Binzel, 1992). Some of the reflected light must be from the coma because 3I/ATLAS is weakly active. This is corroborated by the fact that photometry of $g^{\prime}$- and $r^{\prime}$-band ZTF precovery data (Section 3.1) obtained when the object appeared to be less active, indicated that the $g^{\prime}-r^{\prime}$ colors on those nights were close to solar (within large uncertainties). Specifically, the measurements were $(g^{\prime}-r^{\prime})=0.42\pm 0.14$ on 2025 May 22 and $(g^{\prime}-r^{\prime})=0.44\pm 0.14$ on June 18, compared to $(g^{\prime}-r^{\prime})_{\odot}=0.45\pm 0.02$ (J. Holmberg et al., 2006). This suggests that the much redder color measured from later data could be heavily affected by ejected dust. The spectral slope measured here is slightly higher than the average for cometary dust (see modeling and discussions of typical Solar System comets in S. Protopapa et al. 2018; T. Kareta et al. 2023, as well as Figure 3 of L. Kolokolova et al. 2024), but not significantly so. 2I/Borisov’s coma was similarly red (see, e.g., J. de León et al. 2020), but the object was also significantly more active at the time of the color observations.